what is the values of x so that X^2-2x-15>=0
Answer by checkley77(12569)

X^2-2x-15>=0
(x-5)(x+3)>=0
x-5=0
x=5 ans.
x=3=0
x=-3 ans.Proofs:
5^2-2*5-15>=0
25-10-15>=0
0>=0
-3^2-2*-3-15>=15
9+6-15>=0
0>=0